OpenStudy (anonymous):

Factor the following trinomial: 10x^3 - 71x^2 + 126x

OpenStudy (anonymous):

Factors are: 10: 5*2 & 10*1 126: 126*1, 9*14 & 18*7

OpenStudy (mertsj):

First factor out the common factor which is x

OpenStudy (the_fizicx99):

Try to first factor it as much as possible to make it easier. It doesn't have a coefficient gcf but it does have a variable gcf (they all have x's) Always pick the lowest value, here it's x. x(10x^2 - 71x + 126x) Now multiply 10 * 126, it gives you 1260. Now you find a pair that when multiplied gives you 1260 but adds up to -71. Hint: Use negative numbers

OpenStudy (the_fizicx99):

Opps, x(10x^2 - 71x + 126) *

OpenStudy (the_fizicx99):

Still need help?

OpenStudy (anonymous):

lolololol this guy actually "factored" the trinomial. dam that is too funny.

OpenStudy (anonymous):

Yes, I still need help.

OpenStudy (anonymous):

ok so yes you have to factor out the greatest common factor, so what is the greatest common factor in this equation?

OpenStudy (anonymous):

uhhh, x?

OpenStudy (anonymous):

yep :D so take x and divide it throughout the equation so you get: x(\[10^2-71x+126)\]

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

ok give me one sec.

OpenStudy (the_fizicx99):

There's a shortcut, what number adds 1260 but adds up to -71?

OpenStudy (anonymous):

I have no idea. lol

OpenStudy (whpalmer4):

Try starting with 1/2 of 71. Call it 35. 35 + ??? = 71? Does 35*??? = 1260?

OpenStudy (the_fizicx99):

u-u k... -35 * -36 = 1260 -35 + (-36) = -35 - 36 = -71 10x^2 - 35x - 36x + 126 5x(2x - 7) -18(2x - 7) x[(2x-7)(5x-18)] For some reason I can't use LaTeX, I was going to show you why you move 2x - 7 and 5x - 18 to the other side and still keep them... :/

OpenStudy (whpalmer4):

If not, move to the next number down and try again, until you find a pair that works.

OpenStudy (the_fizicx99):

I prefer to make a list of values.

OpenStudy (whpalmer4):

easy if you have factored it or can do so at sight, less convenient with bigger numbers like 1260. I wouldn't expect anyone asking for help factoring to be able to pluck those numbers out of the air!

OpenStudy (whpalmer4):

factors of 10 that add to -7, sure. Factors of 1260 that add to -71? Maybe 1 in 10 student who has asked me for help factoring would get that without some assistance...

OpenStudy (whpalmer4):

even the factors of 10 that add to -7 is a challenge for quite a few!

OpenStudy (anonymous):

5&2

OpenStudy (the_fizicx99):

Actually I use a much simple method, 1 + 2 + 6 + 0 = 9, 1260/9 = 140 and I can find the values much faster, or because 1260 ends with a 0, it's divisible by 5. Anything that ends in 0 or 5 its divisible by 5

OpenStudy (anonymous):

This stuff is never this difficult in class lol

OpenStudy (the_fizicx99):

It's not really difficult, if you know mental math it make sit much easier!

OpenStudy (the_fizicx99):

makes it*

OpenStudy (anonymous):

Ok so yep I got x\[(5x-18)(2x-7)\] and it works just as @tHe_FiZiCx99 had

OpenStudy (anonymous):

math is never easy lol

OpenStudy (the_fizicx99):

Just requires patience and dedication :) It's another language :D

OpenStudy (anonymous):

it's beyond another language lol

OpenStudy (anonymous):

very true! :D

OpenStudy (the_fizicx99):

cx

OpenStudy (the_fizicx99):

Rose you forgot to add the "x" x[(2x-7)(5x-18)]

OpenStudy (anonymous):

i tried it add it but the program let the x go above the equation… but it is there only hiding (it doesn't want to be in the spotlight) ;P

OpenStudy (the_fizicx99):

it goes inside the \[\ \] or \(\ \)

OpenStudy (the_fizicx99):

Well that failed, \(\ \[ ~and~ \]\) or \(\ \(\ ~and ~ \) \(\color{red}{\)} \)

OpenStudy (anonymous):

lol, I need help with the next one, I think.

OpenStudy (the_fizicx99):

\[\ * and \(\ U_U

OpenStudy (the_fizicx99):

Sure ask it I guess